OpenStudy (anonymous):

can Someone Solve For X? x2 - 2x - 24 = 0

OpenStudy (agreene):

try factoring

OpenStudy (amriju):

hero's formula..

OpenStudy (anonymous):

x² - 2x - 24 = 0 x² - x = 24 + (- 1)² x² - x = 24 + 1 (x - 1)² = 25 x - 1 = ± 5 = x - 1 - 5, = x - 6 = x - 1 + 5, = x + 4 Answer: (x - 6)(x + 4)

OpenStudy (anonymous):

thank you